The U.S. Geological Survey compiled historical data about Old Faithful Geyser (Yellowstone National Park) from 1870 to 1987. Let x1 be a random variable that represents the time interval (in minutes) between Old Faithful eruptions for the years 1948 to 1952. Based on 9160 observations, the sample mean interval was x1 = 62.8 minutes. Let x2 be a random variable that represents the time interval in minutes between Old Faithful eruptions for the years 1983 to 1987. Based on 25,457 observations, the sample mean time interval was x2 = 72.2 minutes. Historical data suggest that σ1 = 8.84 minutes and σ2 = 11.85 minutes. Let μ1 be the population mean of x1 and let μ2 be the population mean of x2.

(a) Compute a 99% confidence interval for μ1μ2. (Use 2 decimal places.)

lower limit    
upper limit
Answer #1

SOLUTION:

From given data,

\bar{x}_{1} =62.8   \bar{x}_{2} = 72.2
\sigma _{1}=8.84 \sigma _{2} = 11.85
n _{1} = 9160 n _{2} = 25457

(a) Compute a 99% confidence interval for μ1μ2. (Use 2 decimal places.)

99% confidence interval

Confidence interval is 99%

99% = 99/100 = 0.99

\alpha = 1 - Confidence interval = 1-0.99 = 0.01

\alpha/2 = 0.01 / 2

= 0.005

Z\alpha/2 = Z0.005  = 2.33

(\bar{x}_{1}-\bar{x}_{2}) \pm Z\alpha/2 * sqrt(\sigma_{1}^{2}/ n _{1} + \sigma_{2}^{2} / n _{2} )  

(\bar{x}_{1}-\bar{x}_{2}) - Z\alpha/2 * sqrt(\sigma_{1}^{2}/ n _{1} + \sigma_{2}^{2} / n _{2} ​​​​​​​​​​​​​​ ) <  \mu _{1}-\mu _{2} <  (\bar{x}_{1}-\bar{x}_{2}) + Z\alpha/2 * sqrt(\sigma_{1}^{2}/ n _{1} + \sigma_{2}^{2} / n _{2} ​​​​​​​​​​​​​​ )

(62.8 -72.2 ) - 2.33 * sqrt(8.842 / 9160  + 11.852 / 25457) <  \mu _{1}-\mu _{2} < (62.8 -72.2 ) + 2.33 * sqrt(8.842 / 9160  + 11.852 / 25457)

-9.4- 2.33 * 0.11852107 <  \mu _{1}-\mu _{2} < -9.4+ 2.33 * 0.11852107

-9.4- 0.2761540931 <  \mu _{1}-\mu _{2} < -9.4+ 0.2761540931

-9.676154 <  \mu _{1}-\mu _{2} < -9.1238459

- 9.68 <  \mu _{1}-\mu _{2} < - 9.12

lower limit : - 9.68 minutes

upper limit: - 9.12 minutes

Please thumbs-up / vote up this answer if it was helpful. In case of any problem, please comment below. I will surely help. Down-votes are permanent and not notified to us, so we can't help in that case.